OpenStudy (anonymous):

What value of q makes the statement true? –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17

OpenStudy (nikato):

start simplifying the right side what is -3+(-14)?

OpenStudy (gorv):

when plus and minus comes together thn it change to minus +-=+

OpenStudy (gorv):

-3-14 now try it

TheSmartOne (thesmartone):

First distribute -q-14=-3-14 Move -14 to the other side by adding -q-14+14=-3-14+14 -q=-3 Divide by -1 on both sides q=3